If 1/2x+3=2/5, then what does 1/2x+1 equal?

Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16946)


(1/2)x+3 = 2/5
(1/2)x + 1 = (2/5)- 2
(1/2)x + 1 = -8/5

Answer by ewatrrr(24785)


Hi
1/2x + 3 = 2/5
1/2x + 3 -2 = 2/5 -2
1/2x + 3 -2 = 2/5 - 10/5
1/2x + 1 = -8/5

Wish You the Best in your Studies.